NSW Oncall log

Instructions

Log

02/07/22

On call: Michelle * Overnight run: MMG-C ran overnight, with several sectors taken out. The reasons were down optical link, unreachable SCA's. * Around midnight, one buffer filled on C14 (which was not in the run). * Nothing was taken out of the run overnight. No filled buffers. Some sectors had corrupt elink packets.

Enrico will install the VS rod busy module on A-Side.

Jared has test with Sergei MMG A to test a new felix release. Enrico's installation will follow that.

Olga is doing an unspecified test at 12 on sTGC side C.

Liang is doing an unspecified test before 12 on sTGC side C.

Aaron will do pulsers on side C around 15:30 after the combined run

Simone will do tests on MMG A after Enrico's ALTI intervention

01/07/22

On call: Michelle * ROC phase calib for sTGC side A with OPC stability. Device 3 (on sector?) buffers fill quickly.Currently looks stable. Issues on sectors 5 6 3 4. Random SCA's in RIM offline. * Zhongliang asks nothing

ROD Busy: * Remove devices needed to remove. Some issue with buffers (what type?), Sergei and Carlo investigating. * Side A unavailable until Enrico is done installing a new ALTI.

Aaron will take pulser runs on mmg C01 in the morning Possibility of threshold scans in the afternoon. Simone wants trigger tests in the afternoon Henri wants to do trimmer run to look at logs, one sector.

-- aawhite CERNCH - 2022-07-01

Comments

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2022-07-01 - AaronStephenWhite
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Sandbox All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2022 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback